OpenStudy (anonymous):

WILL GIVE MEDAL AND FAN. I need someone to help me solve problems not give the answer please.

OpenStudy (anonymous):

When planning his crops, Farmer Joe knows that his 5 acres can support corn and strawberries. He wants to make $5,500 from his crops. He can make $1,000 per acre of corn (the variable c) and $1,500 per acre of strawberries (the variable s). Which equation below would be a constraint in his system of equations?

OpenStudy (anonymous):

c + s = 5 c − s = 5 1,500c + 1,000s = 5,500 1,500s − 1,000c = 5,500

OpenStudy (anonymous):

I think the answer is A